A cord used in bungee jumping is normally 9.88 m long and has spring constant k = 299 N/m. At the lowest point in a jump, the cord length has doubled. How much work has been done on the cord?
Answer #1

Initial length = 9.88 m

Final length = 2×9.88

= 19.76 m

Elongation ,

x = 19.76 - 9.88= 9.88 m

Work done on cord is given by

= (1/2)× 299 × 9.882

= 14600 J